OpenStudy (anonymous):

HELP WILL GIVE A MEDAL: Central park is a rectangle with an area of 840 acres. It is 1/2 a mile wide. One acre is 43,560 square feet. Two people start at the south-west corner at 1:30 pm and start walking round the park in opposite directions. One walks at 5 mph the other walks at 9 mph. How long until they meet again?. A mile is 5280 feet. Answer: Time from start _____ hours

OpenStudy (anonymous):

Central park has an area of 3,400,000 m² It is 800m wide, so it is 3,400,000 m² / 800m = 4250m long. The perimeter is therefore 2 x (800+4250) or 10100m or 10.1 km. One person walks at 1.6 km/hour the other runs at 16 km/hour so the relative speed = 17.6 km/hour Time = distance / speed = 10.1 / 17.6 = 0.57 hours = 34 minutes. Setting off at 1:30, they would meet at 2:04.

OpenStudy (jonnyvonny):

So, after A LOT of conversions, you get: Now, is the easy part. They are going to walk around the part (the perimeter), at certain rates, until they meet. Their rates are additive (meaning you can just combine them) because they are covering different ground at the same time (basically, add them because, their distances combined will equal the perimeter: